# Revisiting Weight Averaging for Model Merging

### v1

@@ 5.2 Main Results
Table 1 presents the comparison with various model merging baselines with ViT-B-32 backbone. First, it shows that the performance of naive weight averaging is far below the upper bound (Individual and Traditional MTL) and generally performs worse compared to the approaches based on task vector (Task Arithmetic and Ties-Merging). However, we observe that applying the proper low-rank approximation to weight averaging by our method significantly boosts the performance, surpassing the prior art based on task vector (Ties-Merging) up to 10.6%. Considering that Ties-Merging employed carefully hand-designed strategies to reduce conflicts among tasks, the improvement suggests that our method can be more effective in resolving inter-task conflicts by filtering out noisy spectral components of task vectors centered at weight averaging.
@@
Our method outperforms even stronger task arithmetic approaches that involve test-time adaptation. For example, AdaMerging optimizes layer-wise coefficients of task vectors during test time, and AdaMerging++ introduces additional mechanisms to reduce inter-task interference. Without such techniques, our method achieves superior performance. When combined with the test-time adaptation (Eq. (6)), our method exhibits further improvement, closing the gap to traditional multi-task learning within 3%. This result indicates that our method can be effectively combined with various merging techniques within the task arithmetic framework, suggesting that it may further close the performance gap with multi-task learning when integrated with more advanced techniques.
Figure 4: Loss landscape visualization of CART with varying rank $k$ and the scaling coefficient $\lambda$ for ViT-B-32 model. We plot the trajectory of the parameters as we increase $k$ from zero to full rank, whose direction is represented as red arrows. The black circle indicates the weight average $\theta_{\text{avg}}$ and the blue star indicates CART with 8% of the full rank, which is used to report the accuracies in Table 1. CART with other ranks are indicated as black crosses. Trajectories with different $\lambda$ are plotted as red dashed lines.
@@
Table 2 presents the results using a larger backbone (ViT-L-14). While the performance of multi-task learning improves with a larger model, we observe even more pronounced gains across all model merging approaches, suggesting that model merging becomes increasingly effective with larger backbones. Consistent with previous observations, we also notice that our method significantly outperforms all baselines, both with and without test-time adaptation. Notably, applying test-time adaptation to our method (CART++) further narrows the gap to the traditional multi-task learning performance within 1%. These results demonstrate that the effectiveness of our method remains consistent across different backbone sizes.
@@ 5.3 Analysis
In this experiment, we assess the impact of the number of tasks on model merging performance. To this end, we construct a subset consisting of $N$ tasks sampled from eight tasks, and report the average performance obtained by all combinations of subsets i.e., $\binom{8}{N}$, while increasing $N$. Figure 6 illustrates the results.
(a) Performance plots for each task on ViT-B-32.
(b) Performance plots for each task on ViT-L-14.
Figure 5: Performance evaluation of CART and Task Arithmetic across eight distinct tasks with low-rank approximation on their task vectors. The plots depict the accuracy (%) of each task as a function of the preserved rank ratio (%) with two different model sizes. CART consistently outperforms Task Arithmetic at a specific rank ratio (e.g., 8%), demonstrating its effectiveness in the model merging process.
Figure 6: An average performance of merged model with increasing number of tasks. We evaluate the model with all possible combinations at each number and report the mean and standard error.
We observe that methods based on simple averaging and task arithmetic exhibit significant performance degradation as the number of tasks increases. This indicates that the inter-task interference intensifies with the increasing number of tasks. On the other hand, the performance of our method remains relatively consistent across varying numbers of tasks. As discussed in Section 4.2, it suggests that eliminating less significant singular vectors of centered task vectors can effectively reduce task interference, leading to a much more scalable approach for model merging.
@@
Since the performance of our method depends on the low-rank approximation of the task vector, we investigate the impact of rank ($k$ in Eq. (4)) in merging performance. Figure 5 presents the per-task performance with varying ranks on two models, ViT-B-32 and ViT-L-14. Although the overall performance depends largely on the rank of task vectors, we observe that the optimal performance lies around $8\%$ of total dimensionality consistently over different tasks and models. It suggests that our method is relatively robust to hyperparameter tuning, reducing the need for extensive searches to find the optimal rank.
@@ 6 Conclusion
In this study, we have revisited the effectiveness of weight averaging from the perspective of model merging. By conducting a detailed analysis to understand why it performs well, and observing these phenomena in the loss landscape, we confirmed that the centered task vectors induced from weight averaging effectively isolates the core knowledge from interfering noise in their singular vectors. Our proposed method, CART, significantly outperformed previous methodologies by a substantial margin across various classification tasks in 8 vision classification datasets. Particularly, through exhaustive experiments on various combinations of tasks, we demonstrated that our method is scalable with respect to the number of merging tasks. Moreover, since our method can be expressed within the framework of task arithmetic, it allows for incremental application of methodologies in this field. We have shown that CART++, an extension of our method, can achieve performance comparable to traditional multi-task learning.
@@ Acknowledgement
This work was in part supported by the National Research Foundation of Korea (RS-2024-00351212, RS-2024-00436165), and Institute of Information and communications Technology Planning and Evaluation (IITP) grant (RS-2024-00509279, RS-2022-II220926, RS-2022-II220959) funded by the Korean government (MSIT).
Appendix A Implementation Details
We provide the implementation details of our experiments.
Detailed Experimental Settings
The baseline codes and pre-trained weights used in our experiments are available in the repository provided by task arithmetic [12]. This includes fine-tuned checkpoints for the eight vision classification tasks (Cars [14], DTD [4], EuroSAT [9], GTSRB [29], MNIST [15], RESISC45 [2], SUN397 [34], and SVHN [22]) across both ViT-B-32 and ViT-L-14 architectures [5]. We used the CLIP [26] implementations from the OpenCLIP  [^1] https://github.com/mlfoundations/open_clip library, which provides the pretrained weights.
Searching the Scaling Coefficient
@@ Appendix A Implementation Details
In CART, we performed a grid search for the scaling coefficient $\lambda\in(0.05,0.1,0.2,\ldots,1.0)$ and selected $\lambda$ that yielded the best performance, following the task arithmetic settings [12]. In CART++, which employs test-time adaptation to optimize the scaling coefficients, we adopted the settings from AdaMerging [36]. Using the test dataset for each task, we iteratively optimized the scaling coefficients separately until the evaluation accuracy converged to an upper bound. We provide the pseudocode for CART and CART++ in Algorithm 1, where the test-time adaptation of CART++ is described in Algorithm 2.
@@
11:     Call Algorithm 2 to optimize $\{\lambda_{t}^{l}\}$ using $\mathcal{D}_{\text{test}}$triangleright CART++
@@
Appendix B Further Analysis on Loss Landscapes
We provide further analysis on the loss landscapes near the model parameters obtained by CART and task arithmetic.

### v2

@@ 5.2 Main Results
Table 1 presents the comparison with various model merging baselines. First, it shows that the performance of naive weight averaging is far below the upper bound (Individual) and generally performs worse compared to the approaches based on task vector (Task Arithmetic). However, we observe that applying the proper low-rank approximation to weight averaging by our method significantly boosts the performance, surpassing the prior art based on task vector (Consensus TA) up to 9.5% for 8 tasks. Considering that Consensus TA employed carefully hand-designed strategies to reduce conflicts among tasks, the improvement suggests that our method can be more effective in resolving inter-task conflicts. Furthermore, this performance advantage persists consistently across experiments involving 14 and 20 tasks, thereby demonstrating the robustness of our proposed methodology with respect to varying numbers of merged tasks.
@@
Our method outperforms strong task arithmetic approaches, including those employing test-time adaptation such as AdaMerge+TA, which optimizes layer-wise coefficients of task vectors at test time. Even without additional techniques, our method achieves superior performance across all numbers of tasks and backbone sizes. When combined with the test-time adaptation (Eq. (8)), our method exhibits further improvement. This result indicates that our method can be effectively combined with various merging techniques within the task arithmetic framework, suggesting that it may further close the performance gap with individually finetuned models when integrated with more advanced techniques.
@@
Table 1 also presents the results using a larger backbone (ViT-L/14). We observe even more pronounced gains across all model merging approaches, suggesting that model merging becomes increasingly effective with larger backbones. Consistent with previous observations, we also notice that our method significantly outperforms all baselines, both with and without test-time adaptation. Notably, applying test-time adaptation to our method (Adamerging+CART) further narrows the gap to individually fine-tuned models. Specifically, with test-time adaptation, the performance gap for the 8-task setting shrinks to approximately 1%, and for the 14- and 20-task settings, the gap remains within 3%. These results demonstrate that the effectiveness of our method remains consistent across different backbone sizes and various numbers of tasks.
Table 2: Multi-Task Performance on eight NLP Tasks with Merged RoBERTa-base with CART.
Method CoLA SST2 MRPC STSB QQP MNLI QNLI RTE Avg.
Individual 0.6018 0.9404 0.8922 0.9063 0.9141 0.8720 0.9271 0.7906 0.8556
Weight Averaging 0.1396 0.6411 0.6936 0.3184 0.7536 0.4219 0.5870 0.5523 0.5134
Task Arithmetic 0.1878 0.8589 0.7990 0.7403 0.8378 0.5908 0.6967 0.6209 0.6665
Ties-Merging 0.2048 0.8440 0.8113 0.5819 0.8570 0.6465 0.7481 0.4296 0.6404
CART (Ours) 0.3294 0.9243 0.8284 0.7956 0.8225 0.5137 0.7985 0.6498 0.7078
Figure 5: Loss landscape visualization of CART with varying rank $k$ and the scaling coefficient $\lambda$ for ViT-B/32 model. We plot the trajectory of the parameters as we increase $k$ from zero to full rank, whose direction is represented as red arrows. The black circle indicates the weight average $\theta_{\text{avg}}$ and the blue star indicates CART with 8% of the full rank. CART with other ranks are indicated as black crosses. Trajectories with different $\lambda$ are plotted as red dashed lines.
Natural Language Processing
Table 5 compares our results against various model merging baselines in the NLP domain. Similar to our findings in the vision tasks, naive weight averaging performs worse than approaches based on task vectors. However, when we apply our low-rank approximation method to weight averaging, performance improves substantially. This improvement suggests that our approach can be applied to various modalities.
@@ 5.3 Analysis
Figure 6: Performance evaluation of CART and Task Arithmetic across eight vision tasks with low-rank approximation on their task vectors. The plots depict the accuracy (%) of each task as a function of the preserved rank ratio (%) of ViT-B/32. (ViT-L/14’s results are provided Appendix C.3) CART consistently outperforms Task Arithmetic at a specific rank ratio (e.g., 8%), demonstrating its effectiveness in the model merging process.
@@
In this experiment, we assess the impact of the number of tasks $N$ on model merging performance. Given the large value of $\binom{20}{N}$, we sampled multiple subsets of $N$ tasks from twenty vision tasks using a statistically sound approach in Appendix D. This sampling, supported by the central limit theorem, ensures a reliable estimate of average performance across all possible combinations. We computed the average normalized performance over these subsets, relative to fine-tuned models, as $N$ increases. Results are presented in Figure 7.
Figure 7: Normalized average performance versus number of tasks for different merging methods applied to ViT-B/32, evaluated on sampled subsets from 20 vision tasks. Mean and standard error are reported.
We observe that methods such as simple averaging (AVG) and task arithmetic (TA) show significant performance degradation as the number of tasks increases. Similarly, Ties-Merging, despite its sophisticated parameter selection strategy, also experiences a notable decrease in performance. In contrast, our method, CART, consistently achieves superior performance across all evaluated numbers of tasks. Moreover, CART demonstrates stability across various task combinations, as indicated by its consistently smaller error bars compared to other methods for all task counts, with the exception of the four-task case where CART and Ties-Merging exhibit comparable error bars. This robustness with respect to the number of tasks and different task combinations can be attributed to CART’s effective reduction of task interference through the centering of task vectors, as detailed in Section 4.2.
@@
Since the performance of our method depends on the low-rank approximation of the task vector, we investigate the impact of rank ($k$ in Eq. (4)) in merging performance. Figure 6 presents the per-task performance with varying ranks on ViT-B/32. The results with ViT-L/14 is presented in the Appendix C.3 (Figure 10). Although the overall performance depends largely on the rank of task vectors, we observe that the optimal performance lies around $8\%$ of total dimensionality consistently over different tasks and models. It suggests that our method is relatively robust to hyperparameter tuning, reducing the need for extensive searches to find the optimal rank.
Extension to model compression
While our approach primarily focuses on model merging i.e., maintaining a single merged model, we can extend it to a model compression that preserves task-specific parameters. In this case, our method also exhibits advantages in terms of compression efficiency by representing each task vector with a small number of rank-1 matrices. In Appendix C.2, we provide a comparative analysis for compression and show that it achieves competitive performance to strong baselines [14, 47].
@@ 6 Conclusion
In this study, we revisit the effectiveness of weight averaging from the perspective of model merging. Through a comprehensive analysis of singular components and loss landscape, we confirm that the centered task vectors induced from weight averaging effectively reduce task interference than the original task vectors, enhancing the performance of model merging. The proposed method, CART, significantly outperforms previous merging approaches by a substantial margin across various tasks and modalities. Moreover, our method can be easily plugged into any merging frameworks based on task arithmetic. We demonstrate this by incorporating CART with test-time adaptation and show that it achieves performance comparable to individually fine-tuned models.
@@ Acknowledgement
This work was in part supported by the National Research Foundation of Korea (RS-2024-00351212, RS-2024-00436165), Institute of Information and communications Technology Planning and Evaluation (IITP) grant (RS-2024-00509279, RS-2022-II220926, RS-2022-II220959) funded by the Korean government (MSIT), and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City.
Supplementary Material
Appendix A Proof for Theorems
In here, we strcitly discuss the relationship between the Row Space Interference and the Task Interference for 1-layer neural network model.
Let $T>2$ be the number of tasks. For each task $t\in[T]$, we have $\mathcal{D}_{t}=\{(x_{t,i},y_{t,i})\}_{i=1}^{n},x_{t,i},\,y_{t,i}\in\mathbb{R}^{d},$ with the linear model $y_{t,i}=(\theta_{0}+\tau_{t})x_{t,i},\quad\theta_{0},\,\tau_{t}\in\mathbb{R}^{d\times d}.$ Each task-specific matrix admits the SVD with non-zero singular values are bounded by some positive constant $[\alpha,s_{max}].$
We assume that inputs lie close to the subspace spanned by row space of $\tau_{t}$ $V_{t}$:
$x_{t,i}=V_{t}\,a_{t,i}+\epsilon_{t,i},\quad a_{t,i}\in\mathbb{R}^{d},\;\|a_{t,i}\|_{2}\leq c\cdot s_{\text{max}},\;\|\epsilon_{t,i}\|_{2}\leq\eta$
where $c$ is a positive constant and $\eta$ is assumed to be a sufficiently small constant (i.e., $\eta\ll c\cdot s_{\text{max}}$).
Theorem 2.
Let a multi task model defined as $W_{\mathrm{MTL}}=\theta_{0}+\sum_{t=1}^{T}\tau_{t}$ and Task Interference $L:=\sum_{t=1}^{T}\sum_{i=1}^{n}\|W_{\text{MTL}}x_{t,i}-(\theta_{0}+\tau_{t})x_{t,i}\|_{2}^{2}$, then $L$ and Row space Interference I satisfies following bound:
$L\leq n(k_{3}I+T(T-1)k_{4}\eta)^{2}$ (9)
where $k_{3}=s_{\text{max}}^{2}c\cdot\frac{rs_{\text{max}}^{2}}{\alpha^{2}},\quad k_{4}=s_{\text{max}}.$
Proof.
By definition of $L$,
$L=\sum_{t=1}^{T}\sum_{i=1}^{n}\left\|\sum_{s\neq t}\tau_{s}x_{t,i}\right\|_{2}^{2}.$
Since $x_{t,i}=V_{t}a_{t,i}+\epsilon_{t,i}$, we have:
$\|\tau_{s}x_{t,i}\|_{2}$ $\leq\|\tau_{s}V_{t}a_{t,i}\|_{2}+\|\tau_{s}\epsilon_{t,i}\|_{2}$
$\leq s_{\text{max}}\|V_{s}^{T}V_{t}a_{t,i}\|_{2}+s_{\text{max}}\|\epsilon_{t,i}\|_{2}$
$\leq s_{\text{max}}\|V_{s}^{T}V_{t}\|_{2}\|a_{t,i}\|_{2}+s_{\text{max}}\|\epsilon_{t,i}\|_{2}$
$\leq s_{\text{max}}(\|V_{s}^{T}V_{t}\|_{2}cs_{\text{max}}+\eta).$
Thus, the norm of the sum over all $s\neq t$ is:
$\left\|\sum_{s\neq t}\tau_{s}x_{t,i}\right\|_{2}$ $\leq\sum_{s\neq t}\|\tau_{s}x_{t,i}\|_{2}$
$\leq\sum_{s\neq t}s_{\text{max}}(\|V_{s}^{T}V_{t}\|_{2}cs_{\text{max}}+\eta).$
Squaring and summing over all $t$ and $i$, we get:
$L$ $\leq\sum_{t=1}^{T}n\left(\sum_{s\neq t}s_{\text{max}}(\|V_{s}^{T}V_{t}\|_{2}cs_{\text{max}}+\eta)\right)^{2}$
$\leq n\left(s_{\text{max}}^{2}c\sum_{t=1}^{T}\sum_{s\neq t}\|V_{s}^{T}V_{t}\|_{2}+T(T-1)s_{\text{max}}\eta\right)^{2}.$
Since $\tau_{t}$ can be composed $U_{t}\Sigma_{t}V_{t}^{T}$ by SVD with $\Sigma_{t}=\operatorname{diag}(s_{t,1},\ldots,s_{t,r_{t}},0,\ldots,0)$ where $r_{t}=rank(\tau_{t})$ and $s_{t,1}\geq s_{t,2}\geq...\geq s_{t,rt}\geq 0$. Also, $\tilde{\Sigma}_{t}=\frac{\Sigma_{t}}{\|\tau_{t}\|_{\text{fro}}}$, where $\|\tau_{t}\|_{\text{fro}}=\sqrt{\sum_{k=1}^{r_{t}}s_{t,k}^{2}}\leq\sqrt{r_{t}}s_{\text{max}}$, by definition of $I$:
$I$ $=\sum_{t=1}^{T}\sum_{s\neq t}\|V_{t}^{T}\tilde{\Sigma}_{t}\tilde{\Sigma}_{s}V_{s}\|_{\text{fro}}$
$\geq\sum_{t=1}^{T}\sum_{s\neq t}\frac{\alpha^{2}}{\|\tau_{t}\|_{\text{fro}}\|\tau_{s}\|_{\text{fro}}}\|V_{t}^{T}V_{s}\|_{2}$
$\geq\sum_{t=1}^{T}\sum_{s\neq t}\frac{\alpha^{2}}{r_{t}s_{\text{max}}^{2}}\|V_{t}^{T}V_{s}\|_{2}.$
Since $r_{t}$ can be bounded by a counstant $r\leq d$, we can bound:
$I\geq\frac{\alpha^{2}}{rs_{\text{max}}^{2}}\sum_{t=1}^{T}\sum_{s\neq t}\|V_{t}^{T}V_{s}\|_{2}.$
Thus,
$\sum_{t=1}^{T}\sum_{s\neq t}\|V_{t}^{T}V_{s}\|_{2}\leq\frac{rs_{\text{max}}^{2}}{\alpha^{2}}I.$
Substituting back into the bound for $L$:
$L$ $\leq n\left(s_{\text{max}}^{2}c\cdot\frac{rs_{\text{max}}^{2}}{\alpha^{2}}I+T(T-1)s_{\text{max}}\eta\right)^{2}$
$\leq n\left(k_{3}I+T(T-1)k_{4}\eta\right)^{2},$
where $k_{3}=s_{\text{max}}^{2}c\cdot\frac{rs_{\text{max}}^{2}}{\alpha^{2}}$, $k_{4}=s_{\text{max}}$ and $r$ is max of $\{rank(\tau_{i})\}_{i=1}^{T}$. ∎
Appendix B Implementation Details
Searching the Scaling Coefficient and Rank
@@ Appendix B Implementation Details
In CART, we performed a grid search for the scaling coefficient $\lambda\in(1.0,0.2,0.4,\ldots,3.0)$ and selected $\lambda$ that yielded the best performance, following the task arithmetic settings [17]. Also, The rank grid can be conducted within the range of $\{4\%,8\%,16\%,32\%\}$, considering different tasks and modalities. In CART++, which employs test-time adaptation to optimize the scaling coefficients, we adopted the settings from AdaMerging [52]. Using the test dataset for each task, we iteratively optimized the scaling coefficients separately until the evaluation accuracy converged to an upper bound. We provide the pseudocode for CART and Adamerging+CART in Algorithm 1, where the test-time adaptation of Adamerging+CART is described in Algorithm 2.
@@
11:      Call Algorithm 2 to optimize $\{\lambda_{t}^{l}\}$ using $\mathcal{D}_{\text{test}}$triangleright Adamerging+CART
@@
Appendix C Additional Results
C.1 Further analysis on I
As depicted in Figure 8, our analysis demonstrates that centered task vector consistently shows low $I(k)$ across all layers than the original task vectors across all preserved rank ratios.
Figure 8: Scatter plot of Row Space interference $I(k)$ for each rank ratio across all layers, with blue and orange points representing Original Task Vector and Centered Task Vector, respectively. Point opacity and size reflect proximity to the median, while solid lines indicate the layer-wise median $I(k)$ for each method.
C.2 Compression Performance
Table 3: Averaged absolute accuracy results on 8, 14 and 20 vision tasks model compresssion benchmark. For CART-Indexing, we retain 8% rank.
Method ViT-B/32 ViT-L/14
8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks
Pretrained $48.0$ $57.1$ $56.0$ $65.0$ $68.4$ $65.4$
Individual $90.5$ $89.6$ $90.5$ $94.2$ $93.3$ $94.1$
With Additional Per-task Parameters
EMR-Merging $88.8$ $85.7$ $86.3$ $93.5$ $91.7$ $92.0$
TALL Mask + Ties $\mathbf{90.8}$ $\mathbf{89.6}$ $\mathbf{90.2}$ $\mathbf{94.3}$ $92.4$ $93.1$
CART-Indexing (Ours) $90.1$ $89.2$ $90.1$ $93.9$ $\mathbf{93.1}$ $\mathbf{93.9}$
Table 4: Additional result of eight NLP tasks with model compression benchmark. For CART-Indexing, we retain 8% rank.
Method CoLA SST2 MRPC STSB QQP MNLI QNLI RTE Avg.
Individual 0.6018 0.9404 0.8922 0.9063 0.9141 0.8720 0.9271 0.7906 0.8556
EMR-Merging 0.3996 0.9335 0.8627 0.8277 0.8972 0.8545 0.8957 0.7437 0.8018
CART-Indexing (Ours) 0.5832 0.9415 0.8897 0.9068 0.9001 0.8721 0.9249 0.7762 0.8493
If only minimal additional per-task parameters are allowed, CART is even further applicable to compression‐based methods thanks to the low‐rank nature of our task vector. Specifically, for each task $t$, we can perform inference as follows:
$\tilde{\mathcal{A}(t)}_{k}=\theta_{\text{avg}}^{l}+\text{SVD}_{k}\bigl{(}\theta_{t}^{l}-\theta_{\text{avg}}^{l}\bigr{)},\quad\forall~{}l\leq L,$
which we refer to as CART-Indexing. As demonstrated in Table 3 and Table 4, our methodology exhibits performance comparable to the sophisticated compression techniques developed by Huang et al. [14] and Wang et al. [47].
For a network comprising $L$ layers, each characterized by an $M\times M$ weight matrix, and tasked with handling $T$ tasks, the binary mask approach requires $T\times L\times M^{2}$ bits for storage. In contrast, our method, which employs low-rank task vectors of rank $k$, demands approximately $32\times T\times L\times 2Mk$ bits. Specifically, with $k=0.08$, our approach achieves a storage cost that is competitive when $M$ is sufficiently large (e.g., $M\approx 1024$), and it becomes increasingly efficient for larger values of $M$.
C.3 Performance of Large Model
In this section, we provide Figure 10 that demonstrate the optimal rank for merging lies around $8\%$ of the full rank for ViT-L/14 as ViT-B/32 shows. Also, we provide the detailed performance of CART for ViT-L/14 in vision classification benchmarks as Figure 9.
Figure 9: Radar plot of normalized accuracy results on 8, 14 and 20 vision classification tasks model merging benchmark for ViT-L-14.
Figure 10: Performance plots for 8 vision task on ViT-L/14.
Appendix D Sampling methods for Task Number scalability Experiments
To obtain Figure 7, we adopted a sampling strategy, selecting a subset of combinations from the 20 vision tasks for each $N$. This approach was necessitated by the computational infeasibility of evaluating all possible combinations of $N$ tasks out of 20, as the total number of combinations, denoted $\binom{20}{N}$, becomes prohibitively large (e.g., $\binom{20}{8}=125,970$). To circumvent this challenge, we relied on the Central Limit Theorem, which asserts that, given a sufficiently large sample size, the distribution of the sample mean approximates a normal distribution, irrespective of the underlying population distribution. Accordingly, for a specified confidence level of $95\%$, a population standard deviation $\sigma$, and a desired margin of error $\epsilon$ in estimating the mean performance, a sample size of at least
$m=\left\lceil\left(\frac{1.96\cdot\sigma}{\epsilon}\right)^{2}\right\rceil$ (10)
combinations is sufficient to ensure an accurate estimate of the true mean. Here, $\sigma$ represents the standard deviation of the performance metric across all possible combinations, and the value 1.96 corresponds to the critical value $z_{0.025}$ from the standard normal distribution for a 95% confidence level. In practice, since $\sigma$ is typically unknown, we employed the Popoviciu inequality to estimate the $\sigma$, which provides an upper bound on the variance of a bounded random variable. For a random variable $X$ constrained to the interval $[a,b]$, the Popoviciu inequality asserts that the variance $\text{Var}(X)$ satisfies:
$\text{Var}(X)\leq\frac{(b-a)^{2}}{4}$
This implies that the standard deviation $\sigma$ is bounded by:
$\sigma\leq\frac{b-a}{2}$ (11)
In our experimental setup, we applied this inequality to establish a upper bound for $\sigma$ for each value of $N$. Specifically, we defined the lower bound $a$ as the average of the bottom $N$ zero-shot performance values across the 20 tasks, reflecting the lowest-performing baseline configurations. Similarly, we set the upper bound $b$ as the average of the top $N$ fine-tuned model performance values, capturing the highest-performing baseline configurations. By determining $a$ and $b$ in this manner, we tailored the upper bound of $\sigma$ to each $N$, ensuring that the sample size $m$ is sufficiently large so that the sample mean reliably estimates the population mean.
To determine hyperparameters for each merge method, we used the default values or recommended hyperparameters presented in each merge method. Specifically, for Task Arithmetic, as defined in Equation 2, we performed a grid search for the merge coefficient $\lambda$ over the range $[0.1,0.2,0.3]$ to identify the optimal value for each task combination. For Ties-Merge, we performed a grid search for the parameter $k$ over $[5,10,20]$, where $k$ represents the percentage of top magnitude parameters to retain in the task vectors, setting the remaining parameters to 0, and fixed the merge coefficient $\lambda=1.0$. For our proposed method, we performed a grid search for the low rank ratio over $[0.08,0.12,0.16]$, and fixed the merge coefficient $\lambda=1.0$.
Appendix E Further Analysis on Loss Landscapes